# Randomness in atomic disorder and consequent squandering of spin-polarization in a ferromagnetically fragile quaternary Heusler alloy FeRuCrSi

### v1

@@ III.5 Magnetic properties
The isothermal magnetization measurements were also carried out below (4 K) and above (300 K) the Curie temperature (Fig. 8). The magnetic isotherm at 4 K exhibit a soft ferromagnetic character with the saturation magnetic moment as 0.97 ${\mu_{B}}$/f.u., much smaller than 2 ${\mu_{B}}$/f.u., expected from the S-P rule described earlier (Sec. III.1). Interestingly, although the Curie temperature in our system is 217 K, the magnetic isotherm at a much higher temperature, 300 K, exhibit a small but definite nonlinear character, suggesting possible presence of ferromagnetic interaction at such high temperature above TC. The presence of ferromagentic interaction, extending upto 300 K and even beyond, has also cast its signature in the magnetic susceptibility results as the inverse susceptibility deviates from linearity below 330 K (inset of Fig. 7). The linear character of inverse susceptibility is the hallmark of the paramagnetic system as suggested by Curie-Weiss law, and its deviation from linearity below 330 K suggests the presence of ferromagnetic clusters persisting in the system at much higher temperature in comparison to its Curie temperature. It may be noted that in literature, a range of Curie temperature (TC) and saturation magnetic moment (Msat), from 346 K to 370 K, and 1.27 ${\mu_{B}}$/f.u. to 1.84 ${\mu_{B}}$/f.u., respectively, have been reported by various groups while investigating the same composition [40, 33]. It is thus quite possible that variation of local disorder in those reported systems are responsible for such variation of TC and Msat. In our case, although the bulk TC is $\sim$217 K, it is thus likely that random distribution of atomic disorder results some local regions having suitable conditions for higher TC in isolated clusters.
III.6 Mössbauer spectrometry
Figure 9: Mössbauer spectra and the multi-component description (see text) of FeRuCrSi taken at 77 K (up) and 300 K (bottom).
Figure 10: Relative fraction as a function of hyperfine field for each distribution in disordered structure of FeRuCrSi. The description excludes the configuration where all the 4c and 4d positions are occupied by Ru atoms only.
Table 3: Refined values of hyperfine parameters for Mössbauer spectra measured at 77 K and 300 K in the FeRuCrSi Heusler alloy. Isomer shift ($\delta$) (quoted relative to $\alpha$-Fe at 300 K), quadrupolar splitting ($\Delta$), quadrupolar shift (${2\varepsilon}$), hyperfine field (Bhf) and relative absorption areas (RA).
T (K) $\delta$ (mm/s) $\Delta$ or ${2\varepsilon}$ (mm/s) Bhf (T) Distribution RA
@@ III.6 Mössbauer spectrometry
To verify the existence of structural disorder and its influence on the magnetic properties of FeRuCrSi, 57Fe Mössbauer spectrometry was carried at 77 K and 300 K. The Mössbauer spectra for both temperatures reveal magnetic sextets superimposed on a central quadrupolar feature. The magnetic contribution very prominent in the spectrum at 77 K, whereas it is much weaker at 300 K. Given that our complete XRD and EXAFS analyses validate a 50:50 site distribution between Fe and Ru at the 4c and 4d positions, one would therefore expect the Mössbauer spectra only two magnetic sextets (associated with positions 4c and 4d) well below the Curie temperature. However, the Mössbauer spectrum taken at 77 K cannot be adequately described with a simple 2 components model. To account physically for this complex hyperfine structure, a minimum of 8 components is required for a satisfactory fit, in agreement with the number of possible environments. The Mössbauer spectrum at 77 K (Fig. 9) is fitted with 7 magnetic sextets, all showing almost similar isomer shift values, and a quadrupolar component (see Table 3), the line widths being fitted with the same value. The small variation of isomer shift could be taken as an indication of variations in the electron density at the Fe nuclei as a function of their atomic environments. The slightly larger isomer-shift align with the various possibilities of atomic environments sharing the same number for each species. In addition, the isomer shift follows a more or less regular increase as the hyperfine field decreases, which includes the quadrupolar component. Another possible description which can be obtained by considering a linear decrease of isomer shift correlated to decreasing hyperfine field values, leads to fairly similar conclusions.
@@
The refined values of the hyperfine parameters are given in Table 3. The average hyperfine field is thus estimated at 16.8 Tesla for such an 8-component fit comprising 7 different magnetic sextets. For the calibration of magnetic moment from hyperfine field values, usually $\alpha$-Fe with a magnetic moment of 2 ${\mu_{B}}$ producing a hyperfine field of 33 Tesla is considered as standard. So, for a 16.8 Tesla magnetic field, a moment of 1.0 ${\mu_{B}}$ is obtained, which matches very closely with the moment obtained from isothermal magnetization Sec. III.5). This is a very surprising result which states that Fe atoms are the primary source to the magnetic moment : indeed, this convincing result is completely inconsistent with all the DFT calculations reported [32, 33], as well as with the one presented in this work (Sec. III.1) which predicts that Cr should be the major source to the magnetic moment, while the magnetic moment at the Fe site would be induced in nature. This is also one of the main reasons why the isothermal saturation magnetization values for this compound deviate from those proposed by the DFT calculations. This also explain why with increasing Fe concentration, the saturation magnetic moment and TC in Ru2-xFexCrSi have been reported to increase linearly [40] and not remain constant to 2 ${\mu_{B}}$/f.u. Fe being the major contributor of magnetic moment in this series also explain why this compound experimentally deviates from the theoretically predicted HMF character.
@@
It is interesting to note, as discussed above, tha the 57Fe Mössbauer spectra could only be analysed by considering 8 different components corresponding to 8 different Fe environments, whereas the XRD and EXAFS measurements suggest that Fe atoms are randomly distributed along Ru atoms only in two crystallographic sites, i.e. 4c and 4d. This apparent contradiction can be easily understood if we consider that the random distribution of Fe atoms resulting from a large number of different localised atomic environments. When Fe is randomly distributed between 4c and 4d sites, there can be 9 different possibilities of atomic distributions, ranging from all 8 atoms being Fe in some unit cells, all 8 Ru in other unit cells, and the rest of the unit cells has varying degree of mixing with the restriction that when averaged over all unit cells, the composition remains FeRuCrSi. These 9 possibilities are: 8 Fe + 0 Ru, 7 Fe + 1 Ru, 6 Fe + 2 Ru, 5 Fe + 3 Ru, 4 Fe + 4 Ru, 3 Fe + 5 Ru, 2 Fe + 6 Ru, 7 Fe + 1 Ru, 0 Fe + 8 Ru. Since 57Fe Mössbauer measurements can only detect Fe atoms, the configuration involving 0 Fe + 8 Ru would not be related to the Mössbauer results, but would provide a very useful nature about the 8 remaining configurations.
@@
As it has been previously mentioned that Fe2CrSi exhibits a ferromagnetic ordering at 520 K [35], while all other members of the Ru-diluted Ru2-xFexCrSi series possess a lower Curie temperature [30, 30, 41], an attempt can be made to correlate the reduction of TC in this series with a reduced magnetic interaction strength caused by non-magnetic (Ru) dilution of the local atomic environment of Fe. Using the same logic in our system, we correlated the different hyperfine magnetic field values of the 8 components of the Mössbauer spectra at 77 K with 8 different atomic environments of the Fe atoms. The largest hyperfine field is shown by the Fe atoms in the unit cell that has only Fe atoms in all the 4c and 4d positions, while the single quadrupolar component which shows no hyperfine field has only 1 Fe atom together with 7 Ru atoms in the unit cell. Excluding the 0 Fe + 8 Ru configuration, our Mössbauer analysis suggests that all the other 8 configurations are present in our system with almost equal probabilities (Table 3). The average value of the magnetic hyperfine field represents the average magnetic interaction which in turn determines the bulk saturation moment as well as TC. This is also corroborated from our Mössbauer spectrum taken at 300 K, where the first four components presented in Table 3 still show weak hyperfine fields, although the bulk TC of our sample, as determined from magnetic susceptibility measurements, is much lower ($\sim$217 K). The non-linearity of the inverse susceptibility (Fig. 7) below $\sim$330 K is probably caused by the presence of such magnetic clusters in our system. The different distribution of the local atomic environment may also explain the variation in Msat and TC reported by various groups for the same equiatomic quaternary Heusler alloy FeRuCrSi. The presence of larger fractions of configurations with more Fe magnetic moments would lead to larger Msat and TC and vice versa. Our results therefore establish a rather rare and complex scenario in which the main physical properties are strongly correlated to the material chemistry in the form of random atomic disorder on a localised scale. This level of understanding of atomic arrangement using Mössbauer spectrometry is remarkable and echoes previous studies based on another related technique, namely nuclear magnetic resonance measurements in other Heusler alloys [12, 54, 55].
@@ III.7 Conclusion
In conclusion, in this work we have used X-ray diffraction, extended X-ray absorption fine structure, 57Fe Mösbauer spectrometry and magnetic measurements in a quaternary Heusler alloy, FeRuCrSi, belonging to Ru2-xFexCrSi series. Since DFT calculations indicate that Cr is a moment carrying atom, all theoretical calculations predict a saturation magnetic moment of 2 $\mu_{B}$/f.u. with robust half-metallicity with 100% spin polarisation for all Fe containing members of this series. However, even for the same composition, there are reports with large variation in Curie temperatures and saturation moments which often deviate significantly from theoretical predictions. Based on our comprehensive structural analysis, we have discovered that the magnetic moment in this system actually comes from Fe atoms, not Cr. The dominant role of Fe in the magnetism explains the experimental deviations from the theoretical predictions and highlights the influence of local atomic disorder on the physical properties. This study highlights the complex relationship between the chemistry of the material, atomic disorder, and the physical properties of the Ru2-xFexCrSi series.

### v2

@@ III.4 Magnetic properties
The isothermal magnetization measurements were also carried out below (4 K) and above (300 K) the Curie temperature of the material (Fig. 7). The magnetic isotherm at 4 K exhibits a soft ferromagnetic character with the saturation magnetic moment as 0.97 ${\mu_{B}}$/f.u., much smaller than 2 ${\mu_{B}}$/f.u., expected from the S-P rule described earlier (Sec. III.1). Interestingly, although the Curie temperature in our system is 216 K, the magnetic isotherm at a much higher temperature, 300 K, exhibits a small but definite nonlinear character, suggesting possible presence of ferromagnetic interaction at such high temperature above $T_{\rm C}$. The presence of ferromagnetic interaction, extending upto 300 K and even beyond, has also cast its signature in the magnetic susceptibility results as the inverse susceptibility deviates from linearity between 216-330 K (inset of Fig. 6). The linear character of inverse susceptibility is the hallmark of the paramagnetic system as suggested by Curie-Weiss law, and its deviation from linearity below 330 K suggests the presence of ferromagnetic clusters persisting in the system at much higher temperature in comparison to its Curie temperature. It may be noted that in the literature, a range of Curie temperature ($T_{\rm C}$) and saturation magnetic moment ($M_{sat}$), from 346 K to 370 K, and 1.27 ${\mu_{B}}$/f.u. to 1.84 ${\mu_{B}}$/f.u., respectively, have been reported by various groups while investigating the same composition [43, 36]. It is thus quite possible that variation of local disorder in those reported systems are responsible for such variation of $T_{\rm C}$ and $M_{sat}$. In our case, although the bulk $T_{\rm C}$ is $\sim$216 K, it is thus likely that random distribution of atomic disorder results in some local regions having suitable conditions for higher $T_{\rm C}$ in isolated clusters.
III.5 Mössbauer spectrometry
Table 3: Refined values of hyperfine parameters for Mössbauer spectra measured at 77 K and 300 K in the FeRuCrSi Heusler alloy: Isomer shift ($\delta$) (quoted relative to $\alpha$-Fe at 300 K), quadrupolar splitting ($\Delta$), quadrupolar shift (${2\varepsilon}$), hyperfine field (Bhf) and relative absorption areas (RA).
$T$ (K) $\delta$ (mm/s) $\Delta$ or ${2\varepsilon}$ (mm/s) $B_{hf}$ (T) Distribution RA
@@ III.5 Mössbauer spectrometry
To verify the existence of structural disorder and its influence on the magnetic properties of FeRuCrSi, 57Fe Mössbauer spectrometry was carried at different temperatures in the range 77 -405 K. The Mössbauer spectra for all temperatures (T $<$ 405 K) reveal multiple magnetic sextets superimposed on a central quadrupolar feature. Fig. 8 clearly shows that the magnetic contribution is very prominent in the spectrum at 77 K, whereas it is much weaker at 300 K. Given that our complete XRD and EXAFS analyses validate a volume average of 50:50 site distribution between Fe and Ru at the 4c and 4d positions, one would therefore expect the Mössbauer spectra to display only two magnetic sextets (associated with positions 4c and 4d) well below the Curie temperature. However, the Mössbauer spectrum taken at 77 K cannot be adequately described with such a simple 2-component model. To account physically for this complex hyperfine structure, a minimum of 8 components is required for a satisfactory fit, in agreement with the number of possible environments. The Mössbauer spectrum at 77 K (Fig. 8) is fitted with 7 magnetic sextets, all showing almost similar isomer shift values, and a quadrupolar component (see Table 3), the line widths being fitted with the same value. Relative fraction as a function of hyperfine field for each distribution in disordered structure of FeRuCrSi is also presented Fig. 9. The small variation of isomer shift could be taken as an indication of variations in the electron density at the Fe nuclei as a function of their atomic environments. The slightly larger isomer-shift align with the various possibilities of atomic environments sharing the same number for each species. In addition, the isomer shift follows a more or less regular increase as the hyperfine field decreases, which includes the quadrupolar component. Another possible description which can be obtained by considering a linear decrease of isomer shift correlated to decreasing hyperfine field values, leads to fairly similar conclusions.
Figure 8: Mössbauer spectra and the multi-component description (see text) of FeRuCrSi taken at 77 K (top), 300 K (middle) and 405 K (bottom). Inset shows zoomed view of the 405 K spectra near origin.
Figure 9: Relative fraction as a function of hyperfine field for each distribution in disordered structure of FeRuCrSi at 77 K. The description excludes the configuration where all the 4c and 4d positions are occupied by Ru atoms only.
@@
The refined values of the hyperfine parameters are given in Table 3. The average hyperfine field is thus estimated at 16.8 Tesla for such an 8-component fit comprising 7 different magnetic sextets. For the calibration of magnetic moment from hyperfine field values, usually $\alpha$-Fe with a magnetic moment of 2 ${\mu_{B}}$ producing a hyperfine field of 33 Tesla is considered as standard. So, for a 16.8 T magnetic field, a moment of 1.0 ${\mu_{B}}$ is obtained, which matches very closely with the moment obtained from isothermal magnetization (Sec. III.4). This is a very surprising result that states that Fe atoms are the primary source to the magnetic moment : indeed, this convincing result is completely inconsistent with all the reported DFT calculations [35, 36], as well as with the one presented in this work (Sec. III.1), which predicts that Cr should be the major source to the magnetic moment, while the magnetic moment at the Fe site would be induced in nature. This is also one of the main reasons why the isothermal saturation magnetization values for this compound deviate from those proposed by the DFT calculations. This also explains why with increasing Fe concentration, the saturation magnetic moment and $T_{\rm C}$ in FexRu2-xCrSi have been reported to increase linearly [43] and not remain constant to 2 ${\mu_{B}}$/f.u. Fe being the major contributor of magnetic moment in this series also explain why this compound experimentally deviates from the theoretically predicted HMF character.
Figure 10: Evolution, as a function of temperature, of the mean values of the hyperfine field estimated from the Mössbauer spectra on the FeRuCrSi powder sample. The inset shows the evolution of the hyperfine field corresponding to the different individual magnetic components. Solid lines are guide to the eye.
@@
It is interesting to note, as discussed above, that the 57Fe Mössbauer spectra could only be analysed by considering 8 different components corresponding to 8 different Fe environments, whereas the XRD and EXAFS measurements suggest that Fe atoms are randomly distributed along Ru atoms only in two crystallographic sites, i.e. 4c and 4d. This apparent contradiction can be easily understood if we consider that the random distribution of Fe atoms resulting from a large number of different localised atomic environments. When Fe is randomly distributed between 4c and 4d sites, there can be 9 different possibilities of atomic distributions, ranging from all 8 atoms being Fe in some unit cells, all 8 Ru in other unit cells, and the rest of the unit cells has varying degree of mixing with the restriction that when averaged over all unit cells, the composition remains FeRuCrSi. These 9 possibilities are: 8 Fe + 0 Ru, 7 Fe + 1 Ru, 6 Fe + 2 Ru, 5 Fe + 3 Ru, 4 Fe + 4 Ru, 3 Fe + 5 Ru, 2 Fe + 6 Ru, 7 Fe + 1 Ru, 0 Fe + 8 Ru. Since 57Fe Mössbauer measurements can only detect Fe atoms, the configuration involving 0 Fe + 8 Ru would not be related to the Mössbauer results, but would provide a very useful nature about the 8 remaining configurations.
Figure 11: Mössbauer spectra at 300K obtained on the FeRuCrSi powder sample without magnetic field (top) and with an external magnetic field of approximately 0.06T (bottom), oriented perpendicular to the $\gamma$-beam). The red and blue lines correspond to the paramagnetic and magnetic components, respectively.
@@
As it has been previously mentioned that Fe2CrSi exhibits a ferromagnetic ordering at 520 K [38], while all other members of the Ru-diluted FexRu2-xCrSi series possess a lower Curie temperature [33, 33, 58], an attempt can be made to correlate the reduction of $T_{\rm C}$ in this series with a reduced magnetic interaction strength caused by non-magnetic (Ru) dilution of the local atomic environment of Fe. Using the same logic in our system, we correlated the different hyperfine magnetic field values of the 8 components of the Mössbauer spectra at 77 K with 8 different atomic environments of the Fe atoms. The largest hyperfine field is shown by the Fe atoms in the unit cell that has only Fe atoms in all the 4c and 4d positions, while the single quadrupolar component which shows no hyperfine field has only 1 Fe atom together with 7 Ru atoms in the unit cell. Excluding the 0 Fe + 8 Ru configuration, our Mössbauer analysis suggests that all the other 8 configurations are present in our system with almost equal probabilities (Table 3 and Fig. 9). The average value of the magnetic hyperfine field represents the average magnetic interaction that in turn determines the bulk saturation moment as well as $T_{\rm C}$. This is also corroborated from our Mössbauer spectrum taken at 300 K, where the first four components presented in Table 3 and inset of Fig. 10 still show weak hyperfine fields, although the bulk $T_{C}$ of our sample, as determined from magnetic susceptibility measurements, is much lower ($\sim 216~{}\mathrm{K}$). A series of Mössbauer spectra were recorded at different temperatures from 77 K until a pure paramagnetic spectrum was obtained at 405 K, as observed in Fig. 8. It is concluded that the Mössbauer spectra recorded in the 150 K–400 K range result from the presence of magnetic sextets and a central quadrupolar component. They were fitted using some magnetic components and a quadrupolar doublet, but the values of some hyperfine parameters have been fixed during the fitting procedure to compensate for the lack of resolution of the central part, illustrated by the broadening of the central component. The mean value of the characteristic hyperfine field for the whole sample is given in Fig. 10. In addition, the values of the hyperfine field for the different components have also been plotted as a function of temperature (Fig. 10: inset). This description confirms the progressive temperature dependence of the magnetically ordered–paramagnetic transition of the Fe moments, which is fully achieved at 405 K.
To better understand the origin of this phenomenon, we compared the hyperfine structures as obtained at 300 K on the powder sample without magnetic field and with an external magnetic field of approximately $0.06~{}\mathrm{T}$ (as shown in Fig. 11), oriented perpendicular to the direction of the $\gamma$-beam produced by a simple magnet placed far from the source (approximately 20 cm) to avoid the presence of polarization effects. The magnetic component proportion increases under the effect of the external magnetic field, which clearly demonstrates the presence of a progressive blocking of the Fe magnetic moments, i.e., the presence of certain dynamic effects of magnetic clusters. The non-linearity of the inverse susceptibility (Fig. 6: inset) below $\sim$330 K is also caused by the presence of such magnetic clusters in our system. The different distribution of the local atomic environment may also explain the variation in $M_{sat}$ and $T_{\rm C}$ reported by various groups for the same equiatomic quaternary Heusler alloy FeRuCrSi. The presence of larger fractions of configurations with more Fe magnetic moments would lead to larger $M_{sat}$ and $T_{\rm C}$ and vice versa.
Our results therefore establish a rather rare and complex scenario in which the main physical properties are strongly correlated to the material chemistry in the form of random atomic disorder on a localised scale. At this stage, it is difficult, on the basis of all the current data, to specify accurately the nature and size of these local atomic environments that lead to a discrete distribution of hyperfine fields, given the number of possible environments associated with the presence of 4 different atomic species. This level of understanding of atomic arrangement using Mössbauer spectrometry is remarkable and echoes previous studies based on another related technique, namely nuclear magnetic resonance measurements in other Heusler alloys [12, 59, 60].
III.6 Electronic structure calculations for disordered crystal structure
Figure 12: Partial density of states for FeRuCrSi for Type-1 disordered structure between Fe and Ru atoms.
[…]
However, the distribution of magnetic moments on each atom in Type-1 (both ordered and disordered) does not agree with the experimental measurements regarding the magnetic ordering. In fact, the DFT calculations for Type-1 show that Cr is the main contributor to the magnetic moment. This is contradicted by the Mössbauer spectrometry and magnetic measurements, which indicate that Fe should be the main contributor to the magnetic moment. This kind of magnetic ordering was reproduced by a Type-3 configuration, with Cr and Ru in a separate plane from Fe (in both ordered and disordered configurations). However, Type-3 is calculated to be slightly less stable than Type-1 at 0 K, but with Fe having a magnetic moment of approximately 2 $\mu_{B}$, antiparallel to the -1 $\mu_{B}$ carried by Cr. This configuration is relatively more consistent with the observations reported in Fig. 3.
The differences between experimental results and theoretical predictions can be attributed to several factors, as the calculations were performed at 0 K on so-called ideal structures. The disordered Type-3 configuration, influenced by temperature or synthesis conditions, could therefore be favored over Type-1, even though structural measurements do not provide a clear distinction. Given all these experimental and theoretical results, no firm conclusion can be drawn on the HMF state of this compound as well as in the entire FexRu2-xCrSi series.
@@ III.7 Conclusion
As a conclusion, this work utilized X-ray diffraction, extended X-ray absorption fine structure, 57Fe Mösbauer spectrometry and magnetic measurements in a quaternary Heusler alloy, FeRuCrSi, belonging to the FexRu2-xCrSi series. These techniques led us to conclude that FeRuCrSi crystallizes in a Type-1 structure with 50:50 site disorder between the Fe and Ru atoms. DFT calculations show this last structure to be the most stable. However, these calculations also indicate that Cr would be the main contributor to the magnetic moment and predict a saturation magnetic moment of 2 $\mu_{B}$/f.u., with robust half-metallicity (100% spin polarization). These theoretical predictions are contradicted by experimental results, showing that Fe is the main contributor to magnetism (saturation moment of 0.97 $\mu_{B}$/f.u.) and challenging the half-metal character. According to DFT calculations, this discrepancy could have been resolved if FeRuCrSi had crystallized in a different structure, where Cr and Ru are placed in a separate plane from Fe (Type–3). However, this arrangement is slightly less stable than Type-1 at 0 K. Despite identical compositions with a mixing of Ru and Fe, there are reports of significant variations in Curie temperatures and saturation moments, which often deviate from theoretical predictions. The dominant role of Fe in the magnetic properties of the sample explains the experimental deviations from theoretical predictions and highlights the impact of the local atomic disorder on the physical properties. The work presented here highlights the complex inter relationship between the chemistry of the material, atomic disorder, and the physical properties of the FexRu2-xCrSi series.